SW&IT Meeting Minutes 2019-01-21
Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang
* Weekly schedules
* Team calendar
* IssueTrak summary
* GitHub open PR
Report auto-generated every week, Do not edit this table, use entries in smartsheets.
| LR20181231 - Bec, Matthieu
| LR20181231 - Colson, Andrew
| LR20181231 - Edmeades, David
| LR20181231 - Sola, Igor
| LR20181231 - Kubanek, Petr
| LR20181231 - Hooper, Stephen
| LR20181231 - Zhang, Yang
| LR20181231 - Zhang, Xianyu
| LR20190114 - Bec, Matthieu
| LR20190114 - Colson, Andrew
| LR20190114 - Edmeades, David
| LR20190114 - Sola, Igor
| LR20190114 - Kubanek, Petr
| LR20190114 - Hooper, Stephen
| LR20190114 - Zhang, Yang
| LR20190114 - Zhang, Xianyu
News from management
Previous Action Items
- Olga Inquirind Mind in IT
Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.
- allsky images named with timestamp
- BCU f/w upgrade documentation
- Dust monitor - bring back
- AO improvement: Luci kwds
- AO cabinet switch upgrade - follow up
- wifi summit upgrade
- Tucson CR power overhaul plan
- n507/obs upgrade
- abell / jhill ws
- OC, WPC, management
- ovms-tt meeting and IPC iterations
- ao-cravat update with Al
- 4d meeting with Bucki, Julian, clarify priorities
- sciarch meeting - limited realease of new portal for testing, release sometimes in Feb.
- Data Mining System = Engineering Science Archive, Allsky, etc. all combined - work in progress
- OAC fixes
- learned how to optimze code with C++11:
- IT #7518- NaN pixels in LBCR images, will communicate with INAF personnel and determine the next step.
- Disabled MODS status query script and updated the MODS wiki to include the steps on how to do it.
- Cleaning up LBC svn -> github conversion.
- Finished a basic AO-Cravat web interface and backend service setup.
- Met with Al Conrad to discuss the next steps. Al will find the details on how to locate calibration frames and update the status of science operations code.
- Provided Al a demo site for him to evaluate.
- Continue working on some tweakings of the web interface and backend server code.
- Investigating several MongoDB stuffs: how to efficiently perform SQL join type query, and other general setups etc.
- Dust particle monitor
- update wiki with user instructions
- verify telemetry
- remove EPICS-EZCA dependency
- make RPM
- update for ALH
- Instabilities (ripples) might be the result of a combination of jumpy actuators and not fully corrected upstream beams
- Main suspect is the LM heaters controller. Suggestion is to remove it from the telescope and send it to Julian Z. at MPE for troubleshooting in his lab.
- Subwindow debugging: #7473
- The problem was an user error, but now the issue is trying to see how to improve the RTD gui to avoid this.
- Looking at #7425 Filter wheel issue. I understand better the issue now but the original coders did already include many works arounds. Mabye Dave can think of a new one?
- Make the MOS unit luci software operate from a remote machine. i.e. not luci.luci.
- AFC GUI stops after exposition, Still not clear why.
- ESM/ARGOS conversion script.
- Michelle like it and discussed a posible more binocular generalization.
- I think it's going to be very similar to the previous week.
- 4D SW interface discussion
- OVMS+TT and Power Backplane meetings
- cheked few IDL adsec routine
- Book the flights for the Italy trip